# Who You Are Matters: Bridging Topics and Social Roles via LLM-Enhanced Logical Recommendation

### v2

@@ Abstract
Recommender systems filter contents/items valuable to users by inferring preferences from user features and historical behaviors. Mainstream approaches follow the learning-to-rank paradigm, which focus on discovering and modeling item topics (e.g., categories), and capturing user preferences on these topics based on historical interactions. However, this paradigm often neglects the modeling of user characteristics and their social roles, which are logical confounders influencing the correlated interest and user preference transition. To bridge this gap, we introduce the user role identification task and the behavioral logic modeling task that aim to explicitly model user roles and learn the logical relations between item topics and user social roles. We show that it is possible to explicitly solve these tasks through an efficient integration framework of Large Language Model (LLM) and recommendation systems, for which we propose TagCF. On the one hand, TagCF exploits the (Multi-modal) LLM’s world knowledge and logic inference ability to extract realistic tag-based virtual logic graphs that reveal dynamic and expressive knowledge of users, refining our understanding of user behaviors. On the other hand, TagCF presents empirically effective integration modules that take advantage of the extracted tag-logic information, augmenting the recommendation performance. We conduct both online experiments and offline experiments with industrial and public datasets as verification of TagCF’s effectiveness, and we empirically show that the user role modeling strategy is potentially a better choice than the modeling of item topics. Additionally, we provide evidence that the extracted logic graphs are empirically a general and transferable knowledge that can benefit a wide range of recommendation tasks.
@@ 1 Introduction
Recommender systems have become an indispensable tool to mitigate information overload and are commonly employed on various online platforms, from e-commerce to video streaming, assisting users in finding personalized content. Traditional recommendation systems rendle2012bpr ; xue2017deep ; he2020lightgcn typically adhere to the learning-to-rank paradigm, which learns the representation vectors of user and item based on the assumption that “similar users exhibit similar behavior”, where these vectors can be interpreted as latent topic distributions, analogous to those in Latent Dirichlet Allocation (LDA) agarwal2010flda distribution. The cornerstone of this paradigm is the discovery and modeling of the item topics (e.g., categories) and how to capture the user preferences on these topics based on historical interactions.
Figure 1: The toy example of the progress from traditional methods to logic collaborative task.
Although effective, this paradigm neglects the modeling of user roles/characteristics and thus fails to capture the logical relationships between user roles and item types, which potentially restricts the expressiveness of the resulting recommendation model. On one hand, existing solutions may discover item-item correlation candidates that are hard to interpret by item types, but the user’s role and personal characteristics may serve as the confounders, providing meaningful explanations for these correlations. A representative real-life example is the famous diaper-beer correlation shen2002objective where a decent amount of human effort has been engaged to find that the “dads with newborns” are the logical explanation for the co-purchase behavior. On the other hand, interest-based modeling mostly relies on interest relations, while the user-item logic relations (e.g., a certain type of user likes a certain type of item) can be far more interpretable and expressive. Consider the intuitive example in Figure 1, we observe a user who purchases a violin after consuming a headphone. While a statistical model may find the violin-headphone relation insignificant, the user may happen to be a symphonist, where both the symphonist-headphone edge and the symphonist-instrument edge are strong and general logical connections. As we will show in section 3.2, knowing the general logic in the real world and what role a user plays in real life may significantly improve the recommender’s ability to engage in more accurate interest exploration liu2024discovery .
@@
Challenges: 1) Different from item topic modeling gupta2010survey ; ahmadian2022deep ; li2023taggpt , for practical and privacy concerns ge2024trustworthy , user role identification is systematically challenging, for it is irresponsible, inefficient, and likely to be offensive to directly ask users to provide their social roles in many web services. Even if the users are willing to provide this information for mutual benefits, there is no guarantee that the provided features are accurate and comprehensive. 2) In terms of the logic modeling task, there have been some pioneering works that use user-generated hashtags or causal tag discovery methods with the help of human experts spirtes2010automated ; vowels2022d ; wang2023causal . However, these methods do not accommodate the scale of industrial recommender systems. Furthermore, they heavily rely on high-quality but manually designed variables, which restrict the model expressiveness in large scale. Ideally, we would like to achieve an automatic modeling framework that can provide an immersive experience where the user roles and the task-specific logic patterns are modeled without bothering the users.
Solution framework: Fortunately, Multi-modal Large Language Models (MLLMs) and Large Language Models (LLMs) have made significant breakthroughs achiam2023gpt ; zhao2023survey ; ouyang2022training ; wu2023next , demonstrating extensive world knowledge memorization abilities and advanced causal and logical reasoning capabilities wei2022chain ; zheng2023judging , which open the opportunities to reexamine the collaborative filtering framework’s ability to model user roles and user behavioral logic. To this end, we propose a general solution framework TagCF that simultaneously solves the aforementioned tasks and improves the recommendation performance. Specifically, we first design a task-specific tag identification module utilizing a MLLM (i.e., M3 cai2024matryoshka ) to extract related user (role) tags and item (topic) tags for each given item, based on the semantic-rich multi-modal features. Then, starting from the identified set of user tags and item tags, we propose a virtual collaborative logic filtering module that uses another LLM (i.e., Qwen2.5-7B yang2024qwen2 ) to iteratively infer the U2I and I2U logic. To meet the scalability demand of the industrial environment, we propose several techniques including cover set reduction and tag-logic model distillation. As we will discuss in Section 3.2.3, this logic graph presents general behavioral logic that can be transferred to other recommendation tasks.
@@ 2 The TagCF Framework
We present the task formulations of the standard top-N recommendation task, the user role (and item topic) identification task, and the behavioral logic reasoning task in Appendix B.1. The key notations in this paper are listed in Appendix B.2.
@@ 2.1 MLLM-based Item-wise Tag Extraction
For a given item $i\in\mathcal{I}$, we first take the original multi-modal information (e.g., audio, image, and title of videos) and use a multi-modal LLM (MLLM), M3 cai2024matryoshka , to generate an semantic item embedding $E_{i}$ and initial textual features. Then, we use the textual features to construct corresponding prompts $X_{i}^{\mathcal{T}}$ for user role tag extraction and $X_{i}^{\mathcal{C}}$ for item topic tag extraction (with prompt details in Appendix B.3). Given $E_{i}$ as auxiliary information, the given prompt will guide the generation of corresponding tags:
@@
Unrestricted Tags and Cover Set Reduction: A critical challenge of the application of Eq.(1) is the unrestricted open-world generation of tags that may gradually accumulate excessive tag sets, while the tag frequency could be extremely skewed (see Appendix D). To circumvent this problem, we propose a greedy and dynamic version of the min cover set finding algorithm (see Appendix B.4) to automatically find a small subset of expressive tags (i.e., the cover set) that provide sufficient coverage of items and are mutually different in semantics. We denote the resulting cover sets of the two tag types as $\mathcal{T}^{\ast}$ and $\mathcal{C}^{\ast}$. In practice, we find that the cover set has some nice features in stability, generality, efficiency, and expressiveness (see Appendix D.1). All these features add up to the effectiveness of the extracted knowledge.
Computational Bottleneck and Distillation: In practice, another key challenge is the computational cost of the MLLMs, especially when there is a large number of newly uploaded items to process on a daily basis (e.g., videos and news). As a countermeasure, we propose to apply Eq.(1) on a smaller subset (tens of thousands) of newly uploaded items, train efficient distilled models $P_{\theta}(t|i):\mathcal{I}\times\mathcal{T}^{\ast}\rightarrow[0,1]$ and $P_{\theta}(c|i):\mathcal{I}\times\mathcal{C}^{\ast}\rightarrow[0,1]$ based on the sampled data, then use $\theta$ to predict user/item tags for all items (in millions). We provide the algorithmic details of this procedure in Appendix B.4. In section 2.3, we show that $\theta$ may also participate in the recommendation model training, providing better alignment with user interactions.
@@ 2.2 LLM-based Collaborative Logic Filtering
With the daily update of $\mathcal{T}$ and $\mathcal{C}$, we use an LLM (i.e., QWen2.5-7B yang2024qwen2 ) to update and maintain the two graphs $\mathcal{G}^{\text{U2I}}$ and $\mathcal{G}^{\text{I2U}}$. Specifically, we iteratively select the tags that have not been included in the (source nodes of) logic graphs, construct the two logic reasoning prompts (in Appendix B.3), then obtain the I2U logic and U2I logic with:
@@
where $X_{c}$ and $X_{t}$ are input prompts for item tag $c\in\mathcal{C}$ and user tag $t\in\mathcal{T}$, $\mathcal{T}_{c}$ and $\mathcal{C}_{t}$ are generated tags, correspondingly. To keep the tag set update inclusive, we also update the tag sets with $\mathcal{T}\leftarrow\mathcal{T}\cup\mathcal{T}_{c}$ and $\mathcal{C}\leftarrow\mathcal{C}\cup\mathcal{C}_{t}$ on a daily basis. Both Eq.(1) and Eq.(2) use the pretrained model without finetuning in order to keep the intact world knowledge and reasoning ability, and we find the generation sufficiently accurate according to human expert justification (Appendix E.1).
Distill Logic within Cover Sets: As we have mentioned in section 2.1, we can achieve a stable and general inference using the cover sets $\mathcal{T}^{\ast}$ and $\mathcal{C}^{\ast}$. However, Eq.(2) does not guarantee a generation output within the cover sets. As a countermeasure, we learn distilled models $P_{\varphi}(c|t)$ and $P_{\varphi}(t|c)$ on the full tag sets with the LLM-inferred data generated by Eq.(2), then predict the logic connections between the cover sets $\mathcal{T}^{\ast}$ and $\mathcal{C}^{\ast}$, where the predicted scores are used to select top-$b$ target tags for each given input tag. We present the details of this process in Appendix B.5 and denote the resulting graph as $\mathcal{G}^{\text{U2I}\ast}$ and $\mathcal{G}^{\text{I2U}\ast}$. Additionally, as we will verify in Section 3.2.3, these graphs are transfer-friendly as they use tags of general concepts and each logic represents a real-world task-agnostic user behavioral logic, taking advantage of the LLM.
@@ 2.3 Tag-Logic Integration in Recommendation
Note that item tags and user tags emphasize different semantic aspects, one can implement two corresponding integration alternatives with symmetric design and we denote them as TagCF-it (that uses item tags to infer) and TagCF-ut (that uses user tags to infer). Without loss of generality, we introduce TagCF-ut with three effective augmentation methods in the following sections, and provide detailed specifications in Appendix B.6.
@@ 2.3.1 Tag-based Encoder
Item Encoder: For each item $i$, we first obtain user tags $\mathcal{T}_{i}$ and item tags $\mathcal{C}_{i}$ through $\theta$ (or Eq.(1)). Then, the embeddings of all extracted tags $\mathbf{T}_{i}=\{\mathbf{e}_{t}|t\in\mathcal{T}_{i}\}$ (or $\mathbf{C}_{i}=\{\mathbf{e}_{c}|c\in\mathcal{C}_{i}\}$ in TagCF-it) are aggregated through either Mean pooling or an Attention Mechanism zhang2019feature (the latter is adopted in practice), generating the tag-based item encoding $\mathbf{r}_{i}^{(t)}\in\mathbb{R}^{d}$ (or $\mathbf{r}_{i}^{(c)}\in\mathbb{R}^{d}$). These encodings provide semantic information that may augment the standard ID-based item embedding $\mathbf{x}_{i}\in\mathbb{R}^{d}$. We provide the details of our attention operation in Appendix B.6.

### v3

@@ Abstract
Recommender systems filter contents/items valuable to users by inferring preferences from user features and historical behaviors. Mainstream approaches follow the learning-to-rank paradigm, which focuses on discovering and modeling item topics (e.g., categories) and capturing user preferences for these topics based on historical interactions. However, this paradigm often neglects the modeling of user characteristics and their social roles, which are logical confounders influencing the correlated interests and user preference transition. To bridge this gap, we introduce the user role identification task and the behavioral logic modeling task that aim to explicitly model user roles and learn the logical relations between item topics and user social roles. We show that it is possible to explicitly solve these tasks through an efficient integration framework of Large Language Model (LLM) and recommendation systems, for which we propose TagCF. On the one hand, TagCF exploits the (Multi-modal) LLM’s world knowledge and logic inference ability to extract realistic tag-based virtual logic graphs that reveal dynamic and expressive knowledge of users, refining our understanding of user behaviors. On the other hand, TagCF presents empirically effective integration modules that take advantage of the extracted tag-logic information, augmenting the recommendation performance. We conduct both online experiments with an industrial environment and offline experiments on public datasets to verify TagCF’s effectiveness, and we empirically show that the user role modeling strategy is potentially a better choice than the modeling of item topics. Additionally, we provide evidence that the extracted logic graphs are empirically a general and transferable knowledge that can benefit a wide range of recommendation tasks. Our code is available in https://github.com/Code2Q/TagCF.
@@ 1 Introduction
Recommender systems have become an indispensable tool to mitigate information overload and are commonly employed on various online platforms, from e-commerce to video streaming, assisting users in finding personalized content. Traditional recommendation systems rendle2012bpr ; xue2017deep ; he2020lightgcn typically adhere to the learning-to-rank paradigm, which learns the representation vectors of user and item based on the assumption that “similar users exhibit similar behavior”, where these vectors can be interpreted as latent topic distributions, analogous to those in Latent Dirichlet Allocation (LDA) agarwal2010flda distribution. The cornerstone of this paradigm is the discovery and modeling of the item topics (e.g., categories) and how to capture user preferences for these topics based on historical interactions.
Figure 1: The toy example of the progress from traditional methods to tag-logic modeling.
Although effective, this paradigm neglects the modeling of user roles/characteristics and thus fails to capture the logical relationships between user roles and item types, which potentially restricts the expressiveness of the resulting recommendation model. On one hand, existing solutions may discover item-item correlation candidates that are hard to interpret by item types, but the user’s role and personal characteristics may serve as the confounders, providing meaningful explanations for these correlations. A representative real-life example is the famous diaper-beer correlation shen2002objective , where a decent amount of human effort has been engaged to find that the “dads with newborns” are the logical explanation for the co-purchase behavior. On the other hand, interest-based modeling mostly relies on interest relations, while the user-item logic relations (e.g., a certain type of user likes a certain type of item) can be far more interpretable and expressive. Consider the intuitive example in Figure 1, we observe a user who purchases a violin after consuming a headphone. While a statistical model may find the violin-headphone relation insignificant, the user may happen to be a symphonist, where both the symphonist-headphone edge and the symphonist-instrument edge are strong and general logical connections. As we will show in section 4.2, knowing the general logic in the real world and what role a user plays in real life may significantly improve the recommender’s ability to engage in more accurate interest exploration liu2024discovery .
@@
Challenges: 1) Different from item topic modeling gupta2010survey ; ahmadian2022deep ; li2023taggpt , for practical and privacy concerns ge2024trustworthy , user role identification is systematically challenging, for it is irresponsible, inefficient, and likely to be offensive to directly ask users to provide their social roles in many web services. Even if the users are willing to provide this information for mutual benefits, there is no guarantee that the provided features are accurate and comprehensive. 2) In terms of the logic modeling task, there have been some pioneering works that use user-generated hashtags or causal tag discovery methods with the help of human experts spirtes2010automated ; vowels2022d ; wang2023causal . However, these methods do not accommodate the scale of industrial recommender systems. Furthermore, they heavily rely on high-quality but manually designed variables, which restricts the model’s expressiveness in a large scale. Ideally, we would like to achieve an automatic modeling framework that can provide an immersive experience where the user roles and the task-specific logic patterns are modeled without bothering the users.
Solution Framework: Fortunately, Multi-modal Large Language Models (MLLMs) and Large Language Models (LLMs) have made significant breakthroughs achiam2023gpt ; zhao2023survey ; ouyang2022training ; wu2023next , demonstrating extensive world knowledge memorization abilities and advanced causal and logical reasoning capabilities wei2022chain ; zheng2023judging , which open the opportunities to reexamine the collaborative filtering framework’s ability to model user roles and user behavioral logic. To this end, we propose a general solution framework TagCF that simultaneously solves the aforementioned tasks and improves the recommendation performance. Specifically, we first design a task-specific tag identification module utilizing an MLLM (i.e., M3 cai2024matryoshka ) to extract related user (role) tags and item (topic) tags for each given item, based on the semantic-rich multi-modal features. Then, starting from the identified set of user tags and item tags, we propose a virtual collaborative logic filtering module that uses another LLM (i.e., Qwen2.5-7B yang2024qwen2 ) to iteratively infer the U2I and I2U logic. To meet the scalability demand of the industrial environment, we propose several techniques, including cover set reduction and tag-logic model distillation. As we will discuss in Section 4.2.3, this logic graph presents general behavioral logic that can be transferred to other recommendation tasks.
@@
2 Related Work
2.1 Collaborative Filtering
Collaborative filtering (CF), one of the most successful recommendation approaches, continues to attract interest in both academia and industry. Over time, CF has evolved from traditional methods sarwar2001item ; linden2003amazon ; breese2013empirical ; akritas2004applications ; mnih2007probabilistic to advanced techniques incorporating sequences hidasi2015session ; kang2018self ; sun2019bert4rec and graph structures wang2019neural ; he2020lightgcn . Among the representative methods, matrix factorization (MF) techniques akritas2004applications ; mnih2007probabilistic are effective in learning latent user and item representations. Sequential CF methods extend this by modeling the temporal order of user interactions with Recurrent Neural Networks hidasi2015session and Transformers kang2018self ; sun2019bert4rec ; liu2024mamba4rec . Graph-based CF methods like NGCF wang2019neural and LightGCN he2020lightgcn have also gained attention in recent years. Besides, self-supervised learning approaches yu2023xsimgcl ; cai2023lightgcl ; zhang2024recdcl have been explored to enhance CF by learning robust representations. However, these methods often ignore user roles and logical relationships between characteristics.
Meanwhile, some personality-aware filtering methods incorporate user traits through neighborhood filtering khelloufi2020social ; dhelim2020compath ; dhelim2023hybrid or matrix factorization extensions khelloufi2020social ; dhelim2020compath . In the literature of psychology goldberg2013alternative , the majority of the works used the Big-Five personality model to represent the user’s personality, while the choice of the most suitable personality definition that satisfies the requirements of the recommendation application still needs further investigation. Recent works li2023gpt4rec ; yang2023palr ; shang2024personalized have attempted to leverage LLMs for personalized recommendations and user interest interpretation. While progress has been made, existing approaches still overlook explicit modeling of user roles and their logical relationships. In this work, we aim to address these gaps by bridging topics and social roles via LLMs-enhanced logical recommendation within the CF framework.
2.2 LLM-based Recommendation
LLM-enhanced Recommender. Many current works have explored how to apply the LLM to generate auxiliary knowledge for enhancing traditional RS. LLMRG wang2024llmrg fabricates prompts to construct chained graph reasoning from LLM to augment the recommendation model. LLMHG chu2024llm first leverages LLMs to deduce Interest Angles (IAs) and categorize movies into multiple categories within each IA to construct a multi-view hypergraph. SAGCN  liu2023understanding uses a chain-based prompting strategy to extract semantic interactions from LLM for each review and introduces a semantic aspect-based graph convolution network to enhance the user and item representations by leveraging these semantic aspect-aware interactions. LLM-KERec zhao2024breaking uses LLM to identify the complementary relationships of an item knowledge graph. Subsequently, they train an entity-entity-item weight decision model which is then used to inject knowledge into the ranking model by using the real exposure and click feedback of complementary items. Nevertheless, current methods of using LLM for data enhancement primarily focus on the meta-features, neglecting knowledge from the user side and the logic rationale between user-item interactions. This limitation hinders their ability to facilitate traditional recommenders to capture semantic and representative collaborative information.
LLM as Recommender Itself. Recently, LLMs have demonstrated remarkable performance across a wide range of recommendation tasks. P5 geng2022recommendation and M6Rec cui2022m6 finetune LLM by modeling recommendation tasks as natural language processing tasks. ChatRec gao2023chat employs LLMs as a recommender interface for conversational multi-round recommendations. TALLRec bao2023tallrec designs a customized parameter-efficient tuning process for recommendation tasks on LLM with a LoRA architecture. HLLM  chen2024hllm uses an item LLM to encode text features, feeding its embeddings to a user LLM for recommendations. Compared to LLM-enhanced recommender, this paradigm’s computational cost (both for training and inference) is much higher and the industry-deployable solution is still an open question deng2025onerec . As we will discuss in section 3.3.1, this research direction focuses on the improvement of sequential models, which is complementary to our proposed knowledge extraction and augmentation framework.
@@ 3 The TagCF Framework
We present the task formulations of the standard top-N recommendation task, the user role (and item topic) identification task, and the behavioral logic reasoning task in Appendix A.1. The key notations in this paper are listed in Appendix A.2.
@@ 3.1 MLLM-based Item-wise Tag Extraction
For a given item $i\in\mathcal{I}$, we first take the original multi-modal information (e.g., audio, image, and title of videos) and use a multi-modal LLM (MLLM), M3 cai2024matryoshka , to generate a semantic item embedding $E_{i}$ and initial textual features. Then, we use the textual features to construct corresponding prompts $X_{i}^{\mathcal{T}}$ for user role tag extraction and $X_{i}^{\mathcal{C}}$ for item topic tag extraction (with prompt details in Appendix A.3). Given $E_{i}$ as auxiliary information, the given prompt will guide the generation of tags:
@@
Unrestricted Tags and Cover Set Reduction: A critical challenge of the application of Eq.(1) is the unrestricted open-world generation of tags that may gradually accumulate excessive tag sets, while the tag frequency could be extremely skewed (see Appendix C). To circumvent this problem, we propose a greedy and dynamic version of the min cover set finding algorithm (see Appendix A.4) to automatically find a small subset of expressive tags (i.e., the cover set) that provide sufficient coverage of items and are mutually different in semantics. We denote the resulting cover sets of the two tag types as $\mathcal{T}^{\ast}$ and $\mathcal{C}^{\ast}$. In practice, we find that the cover set has some nice features in stability, generality, efficiency, and expressiveness (see Appendix C.1). All these features add up to the effectiveness of the extracted knowledge.
Computational Bottleneck and Distillation: In practice, another key challenge is the computational cost of the MLLMs, especially when there is a large number of newly uploaded items to process on a daily basis (e.g., videos and news). As a countermeasure, we propose to apply Eq.(1) on a smaller subset (tens of thousands) of newly uploaded items, train efficient distilled models $P_{\theta}(t|i):\mathcal{I}\times\mathcal{T}^{\ast}\rightarrow[0,1]$ and $P_{\theta}(c|i):\mathcal{I}\times\mathcal{C}^{\ast}\rightarrow[0,1]$ based on the sampled data, then use $\theta$ to predict user/item tags for all items (in millions). We provide the algorithmic details of this procedure in Appendix A.4. In section 3.3, we show that $\theta$ may also participate in the recommendation model training, providing better alignment with user interactions.
@@ 3.2 LLM-based Collaborative Logic Filtering
With the daily update of $\mathcal{T}$ and $\mathcal{C}$, we use an LLM (i.e., QWen2.5-7B yang2024qwen2 ) to update and maintain the two graphs $\mathcal{G}^{\text{U2I}}$ and $\mathcal{G}^{\text{I2U}}$. Specifically, we iteratively select the tags that have not been included in the (source nodes of) logic graphs, construct the two logic reasoning prompts (in Appendix A.3), then obtain the I2U logic and U2I logic with:
@@
where $X_{c}$ and $X_{t}$ are input prompts for item tag $c\in\mathcal{C}$ and user tag $t\in\mathcal{T}$, $\mathcal{T}_{c}$ and $\mathcal{C}_{t}$ are generated tags, correspondingly. To keep the tag set update inclusive, we also update the tag sets with $\mathcal{T}\leftarrow\mathcal{T}\cup\mathcal{T}_{c}$ and $\mathcal{C}\leftarrow\mathcal{C}\cup\mathcal{C}_{t}$ on a daily basis. Both Eq.(1) and Eq.(2) use the pretrained model without finetuning in order to keep the intact world knowledge and reasoning ability, and we find the generation sufficiently accurate according to human expert justification (Appendix D.1).
Distill Logic within Cover Sets: As we have mentioned in section 3.1, we can achieve a stable and general inference using the cover sets $\mathcal{T}^{\ast}$ and $\mathcal{C}^{\ast}$. However, Eq.(2) does not guarantee a generation output within the cover sets. As a countermeasure, we learn distilled models $P_{\varphi}(c|t)$ and $P_{\varphi}(t|c)$ on the full tag sets with the LLM-inferred data generated by Eq.(2), then predict the logic connections between the cover sets $\mathcal{T}^{\ast}$ and $\mathcal{C}^{\ast}$, where the predicted scores are used to select top-$b$ target tags for each given input tag. We present the details of this process in Appendix A.5 and denote the resulting graph as $\mathcal{G}^{\text{U2I}\ast}$ and $\mathcal{G}^{\text{I2U}\ast}$. Additionally, as we will verify in Section 4.2.3, these graphs are transfer-friendly as they use tags of general concepts and each logic represents a real-world task-agnostic user behavioral logic, taking advantage of the LLM.
@@ 3.3 Tag-Logic Integration in Recommendation
Note that item tags and user tags emphasize different semantic aspects, one can implement two corresponding integration alternatives with symmetric design and we denote them as TagCF-it (that uses item tags to infer) and TagCF-ut (that uses user tags to infer). Without loss of generality, we introduce TagCF-ut with three effective augmentation methods in the following sections, and provide detailed specifications in Appendix A.6.
@@ 3.3.1 Tag-based Encoder
Item Encoder: For each item $i$, we first obtain user tags $\mathcal{T}_{i}$ and item tags $\mathcal{C}_{i}$ through $\theta$ (or Eq.(1)). Then, the embeddings of all extracted tags $\mathbf{T}_{i}=\{\mathbf{e}_{t}|t\in\mathcal{T}_{i}\}$ (or $\mathbf{C}_{i}=\{\mathbf{e}_{c}|c\in\mathcal{C}_{i}\}$ in TagCF-it) are aggregated through either Mean pooling or an Attention Mechanism zhang2019feature (the latter is adopted in practice), generating the tag-based item encoding $\mathbf{r}_{i}^{(t)}\in\mathbb{R}^{d}$ (or $\mathbf{r}_{i}^{(c)}\in\mathbb{R}^{d}$). These encodings provide semantic information that may augment the standard ID-based item embedding $\mathbf{x}_{i}\in\mathbb{R}^{d}$. We provide the details of our attention operation in Appendix A.6.